OpenStudy (anonymous):

Astronomers measure large distances in light-years. One light-year is the distance that light can travel in one year, or approximately 5.88 × 10^12 miles. Suppose a star is 9.8 × 10^1 light-years from Earth. In scientific notation, approximately how many miles is it?

OpenStudy (anonymous):

Each light year is about 5.88e12 miles. So, to find this measurement in miles, we'll take the amount of light years and multiply it by this conversion factor: \[98 light years \times \frac{ 5.88 \times 10^{12} miles }{ 1 light year }=?\]The light years will cancel out and you'll be left with miles.

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

Multiply the star's distance in light years by the number of miles in a light year.

OpenStudy (anonymous):

2a^-4m5 what is the answer if you simplify the expression

OpenStudy (anonymous):

5 is also a exponent

OpenStudy (anonymous):

Where are you getting the "a" and "m" from?

OpenStudy (anonymous):

\[\left(9.8*10^1\right)*\left(5.88*10^{12}\right)=5.7624*10^{14} \text{ miles} \]